# Non-monotonic emergence of order from chaos in turbulent thermo-acoustic fluid systems

### v3

@@ III Results
Further, the variation of CPL and $\overline{C}_{BC}$ with $p^{\prime}_{rms}$ derived from a bluff-body and swirl-stabilized turbulent combustor is shown in Fig. 5(a) and Fig. 5(b), respectively. Clearly, both measures are minimum during chaotic dynamics and reach a maximum during periodic dynamics (thermo-acoustic instability). We infer that the variation in the magnitudes of jumps in amplitude scales due to fluctuations decreases in a monotonic fashion as order emerges amidst chaos in the system. Note that, as $Re$ increases in the regime $Re<Re_{c}$, fluctuations become more random and uncertain, and hence, the network entropy ($\mu_{S}$) increases, while the amplitude jumps associated with the fluctuations become less diverse. For $Re>Re_{c}$, both the number of possible transitions and the range of amplitude scales in which the fluctuations occur decrease. Moreover, CPL and $\overline{C}_{BC}$ exhibit distinct power-law scaling behavior with the root-mean-square of acoustic pressure fluctuations. The power law exponents are nearly identical for experiments in both bluff-body and swirl-stabilized turbulent combustor.
@@
In summary, we used complex networks to characterize the fluctuations and the jumps in amplitudes due to fluctuations as a dynamical transition occurs in a complex turbulent thermo-fluid system. Intriguingly, during the transition from chaos to order, we discover a regime where the amount of uncertainty in fluctuations increases, but the amplitude jumps become less diverse.
IV Conclusion
@@
We study the dynamical transition from chaos to order in a turbulent thermo-fluid by analyzing the fluctuations in one of the subsystems. Fluctuations in system variables reflect as well as cause coupling between sub-processes of a complex system. We use a complex network approach to encode the fluctuations in the system dynamics and apply network measures to quantify the characteristics of fluctuations during the transition from chaos to order in the system. The network is constructed by binning amplitudes and encoding the transition probabilities due to fluctuations between these bins and is referred to as amplitude transition networks. The average entropy $\mu_{S}$ of an amplitude transition network captures the uncertainty in the amount of fluctuations between different amplitude scales and reflects the complexity of the dynamics at a fixed time scale. On the other hand, distance-based network metrics (CPL and $\overline{C}_{BC}$) indicate the diversity in the jumps in amplitude scales due to fluctuations in the time series. Thus, our study offers a novel outlook in characterizing fluctuations, which plays a crucial role in order emergent phenomena [18].
@@
We are able to identify the critical value $Re_{c}$, where the inter-subsystem interactions begin to dominate the influence of the underlying turbulence, allowing self-sustained ordered dynamics to emerge in the flow. The role of fluctuations in a complex system exhibiting emergent dynamics can be analyzed for various natural and engineering systems using the approach presented in this work. By encoding the fluctuations as the topology of a network, we also encode the structure of interactions in the system that influence these fluctuations. Additionally, we propose that using measures derived from amplitude transition networks, one can quantify as well as compare the different types of dynamical transitions across systems, such as monotonic, non-monotonic, or abrupt transitions between chaos and order.
@@ Acknowledgements.
The authors thank S. Sivakumar, S. Thilagaraj, S. Anand (Indian Institute of Technology, Madras), P. R. Midhun (Aalto University, Finland), Anaswara Bhaskaran (University of Pavia, Italy, and, Dr. Samadhan A. Pawar (University of Toronto) for providing the experimental data. This research was supported by J. C. Bose Fellowship (JCB/2018/000034/SSC) and the IoE-IITM Research Initatives (SP/22-23/1222/CPETWOCTSHOC) grants provided to R.I.S.. S.T. acknowledges the support from Prime Minister’s Research Fellowship, Govt. of India.
@@
Appendix A Effect on the number of amplitude bin used in network construction

### v4

@@ III Results
Further, the variation of CPL and $\overline{C}_{BC}$ with $p^{\prime}_{rms}$ derived from a bluff-body and swirl-stabilized turbulent combustor is shown in Fig. 5(a) and Fig. 5(b), respectively. Clearly, both measures are minimum during chaotic dynamics and reach a maximum during periodic dynamics (thermo-acoustic instability). We infer that the variation in the magnitudes of jumps in amplitude scales due to fluctuations decreases in a monotonic fashion as order emerges amidst chaos in the system. Intriguingly, during the transition from chaos to order with increase in $Re$, we discover a regime ($Re<Re_{c}$) where fluctuations become more disordered and uncertain, and hence, the network entropy ($\mu_{S}$) increases, while the amplitude jumps associated with the fluctuations become less diverse. For $Re>Re_{c}$, both the number of possible transitions and the range of amplitude scales in which the fluctuations occur decrease. Moreover, CPL and $\overline{C}_{BC}$ exhibit distinct power-law scaling behavior with the root-mean-square of acoustic pressure fluctuations. The power law exponents are nearly identical for experiments in both bluff-body and swirl-stabilized turbulent combustors.
@@
In summary, we used complex networks to characterize the fluctuations and the jumps in amplitudes due to fluctuations as a dynamical transition occurs in a complex turbulent thermo-fluid system.
IV Discussion and Conclusions
@@
We study the dynamical transition from chaos to order in a turbulent thermo-fluid system by analyzing the fluctuations in one of the subsystems. Fluctuations in system variables reflect the interactions between underlying sub-processes. At the same time, these fluctuations also affect the functional form of the sub-units of a complex system. We use a complex network approach to encode the fluctuations in the system dynamics and apply network measures to quantify the characteristics of fluctuations during the transition from chaos to order in the system. The network is constructed by binning amplitudes and encoding the transition probabilities due to fluctuations between these bins and is referred to as amplitude transition networks. The average entropy $\mu_{S}$ of an amplitude transition network captures the uncertainty in the amount of fluctuations between different amplitude scales and reflects the complexity of the dynamics at a fixed time scale. On the other hand, distance-based network metrics (CPL and $\overline{C}_{BC}$) indicate the diversity in the jumps in amplitude scales due to fluctuations in the time series. Thus, our study offers a novel outlook in characterizing fluctuations, which plays a crucial role in the emergent of ordered dynamics [18].
@@
We identify a critical value $Re_{c}$, such that for $Re<Re_{c}$, disorder increases with $Re$ in the form of uncertainty in fluctuations in the acoustic pressure. However, beyond $Re_{c}$, disorder decreases despite the increase in $Re$ owing to strong positive feedback and organized interactions between subsystems. Self-organization becomes prominent beyond $Re_{c}$ leading to the emergence of self-sustained ordered dynamics in the acoustic field despite the turbulence in the underlying flow and flame dynamics. The role of fluctuations in a complex system exhibiting emergent dynamics can be analyzed for various natural and engineering systems using the approach presented in this work. By encoding the fluctuations as the topology of a network, we also encode the structure of interactions in the system that influence these fluctuations. Additionally, we propose that using measures derived from amplitude transition networks, one can quantify as well as compare the different types of dynamical transitions across systems, such as monotonic, non-monotonic, or abrupt transitions between chaos and order.
An analytical characterization of the non-monotonic transition would explain several features of the dynamical transition in turbulent thermo-acoustic systems would be an interesting challenge for future work. Recently, the transition to periodic oscillations was modeled as a transition from defect to phase turbulence [41]. This was achieved using several Stuart Landau oscillators that are both diffusively and globally coupled through a globally coupled Complex-Ginzburg Landau equation, capturing both qualitative and quantitative features of the system. A possible direction would be extending this model to study the relation between amplitude fluctuations and dynamical transition in the acoustic field.
@@ Acknowledgements.
The authors thank S. Sivakumar, S. Thilagaraj, S. Anand (Indian Institute of Technology, Madras), P. R. Midhun (Aalto University, Finland), Anaswara Bhaskaran (University of Pavia, Italy) and Dr. Samadhan A. Pawar (University of Toronto) for providing the experimental data. This research was supported by J. C. Bose Fellowship (JCB/2018/000034/SSC) and the IoE-IITM Research Initatives (SP/22-23/1222/CPETWOCTSHOC) grants provided to R.I.S.. S.T. acknowledges the support from Prime Minister’s Research Fellowship, Govt. of India. The authors also acknowledge Mr. Radhakrishnan and Dr. Raghunathan for their valuable discussions.
@@
Appendix A Characterizing intermittency dynamics using recurrence plots
The transition from chaos to ordered dynamics via the route of intermittency in turbulent thermo-acoustic systems was reported first by Nair et al. [31] and subsequently by others [64, 65, 66, 67, 68, 69, 70].
Recurrence plots help to visualize and quantify the architectural features of a time series [71]. These plots encode the time taken for a trajectory to recur at a location in the phase space. The recurrence matrix has entries of ones (1s) corresponding to the time stamps when the dynamics recur in the phase space and zeros otherwise. The patterns in the recurrence matrix characterize the dynamical state of the system. For example, parallel diagonal lines signify the periodic nature of the time series, and the distance between the parallel lines denote the time period of oscillations.
Figure 6: (a) Time series of acoustic pressure during intermittency ($Re=(4.854\pm 0.039)\times 10^{4}$) and its corresponding recurrence matrix. The insets (b) and (c) shows the topology of recurrence matrix during periodic and aperiodic fluctuation epochs of the pressure signal, respectively.
Recurrence plots can help to identify and distinguish intermittency dynamics in time series data [72]. A time series is embedded in a phase space using Takens’ delay embedding theorem [73]. According to Takens’ theorem, we construct the phase space using the delayed time series $x(t+n\tau)$, where $n\in[0,m]$ and $m$ is the embedding dimension. Also, $\tau$ is the time lag for which the time series $x(t)$ and $x(t+\tau)$ have minimum mutual information. Next, a recurrence plot is constructed such that both the abscissa and ordinate represent time stamps. If the phase space trajectory visits the same location (within a reasonable bound $\varepsilon$) at two different time stamps $t_{i}$ and $t_{j}$, then the entry in the recurrence plot at $(t_{i},t_{j})$ and $(t_{j},t_{i})$ is set to value one indicating recurrence of the trajectory in the phase space. Further details for phase space construction and recurrence plot construction are presented in Nair et al. [31].
[…]
The recurrence matrix shown in Fig. 6 has regions of parallel diagonal lines (Fig. 6(b)) interspersed between black patches (Fig. 6(c)). This topology of the recurrence matrix depicts the intermittent nature of acoustic pressure signal as there is a co-existence of black patches signifying aperiodicity and diagonal lines signifying periodicity [72]. The black patches correspond to low-amplitude aperiodic fluctuations (see $t\in[2.5,2.6]$; Fig. 6(c)) and the diagonal lines represent the high-amplitude periodic oscillations (see $t\in[2.1,2.2]$; Fig. 6(c)).
Appendix B Effect on the number of amplitude bin used in network construction